A lab assistance want to make 5 liters 25.6% acid solution. if solutions of 40% and 16% are in stock, how many liters of each must be mixed to prepare the solution?

Answer by oberobic(2304)

You need 5 liters total. This needs to be 25.6% 'pure' acid.
.256*5 = 1.28 liters of pure acid in the 5 liters total solution.
.
The amount of the 40% solution is: x
so, the amount of the 16% solution is: 5-x
.
.40x + .16(5-x) = 1.28
.
multiply by 100 to eliminate the decimal
.
40x + 16(5-x) = 128
.
40x + 80 -16x = 128
.
24x = 48
.
x = 2
.
so
.
5-x = 3
.
Check to see if the amount of pure acid is correct.
.
.4*2 + .16*3 = 1.28
which is exactly the amount of pure acid calculated at the start.
.
Answer: Mix 2 liters of 40% acid and 3 liters of 16% acid to make 5 liters of 25.6% acid.
.
Done.